OpenStudy (celinegirl):

Leo bought a bulldozer for $63,103. The value of the bulldozer depreciated at a constant rate per year. The table below shows the value of the bulldozer after the first and second years: Year 1 2 Value (in dollars) 58,054.76 53,410.38

OpenStudy (anonymous):

Looks like one number..can't tell the start

OpenStudy (celinegirl):

sorry my computer was lagging.. let me post the rest of the question

OpenStudy (celinegirl):

Which function best represents the value of the bulldozer after t years? f(t) = 58,054.76(0.92)t f(t) = 63,103(0.08)t f(t) = 63,103(0.92)t f(t) = 58,054.76(0.08)t

OpenStudy (celinegirl):

oops the t's at the end are exponents. and i think its C

OpenStudy (anonymous):

C is the same price as it was originally. Is there anymore to the question?

OpenStudy (celinegirl):

no thats it

OpenStudy (anonymous):

Well two of the answers are the same except for the end part, trying to think what it is.. what course are you in for mathematics?

OpenStudy (celinegirl):

algebra 1

OpenStudy (anonymous):

hm..still stumped. Maybe someone else has an idea..

OpenStudy (celinegirl):

yeah

OpenStudy (retireed):

celinegirl do you have a graphing calculator, like a TI-84 ?

OpenStudy (celinegirl):

no